How far would a car travel in 6.0 s if it started from rest, and accelerated at 2.0m/s2?
Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(35256)

d = v[i]*t + (1/2)*a*t^2

d = (0)(6) + (1/2)*(2)*(6)^2

d = (0)(6) + (1/2)*(2)*(36)

d = 0 + (1/2)*(72)

d = 0 + 36

d = 36

So the car would travel 36 meters

Answer by solver91311(24713)









For the given situation, and , so

meters.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism